# Space-based mm/mg-scale Laser Interferometer for Quantum Gravity

### v1

@@ Appendix A State space model
[…]
@@
[…]
@@
Thus, $\gamma_{\pm}$ represent the full width at half-maximum (FWHM). To demonstrate the GIE, the condition $\gamma_{+}/2>n_{+}\Gamma$ represents an approximate threshold, as depicted by the red curve in Fig. 2 of the main text.
Appendix B Covariance matrix
@@
$\bm{\tilde{V}}_{\pm}=\left(\begin{array}[]{cc}V^{\pm}_{qq}&V^{\pm}_{qp}\\ V^{\pm}_{qp}&V^{\pm}_{pp}\end{array}\right).$ (B21)
@@
$\bm{\mathcal{V}}\equiv\Bigl{(}\begin{array}[]{cc}\bm{\mathcal{V}}_{1}&\bm{\mathcal{V}}_{12}\\ \bm{\mathcal{V}}_{12}&\bm{\mathcal{V}}_{2}\end{array}\Bigr{)}=\bm{S}\Bigl{(}\begin{array}[]{cc}\bm{\tilde{V}}_{+}&0\\ 0&\bm{\tilde{V}}_{-}\end{array}\Bigr{)}\bm{S}^{\text{T}},~{}~{}$ (B26)
$\bm{S}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cccc}1&0&1/(1-\delta)^{1/4}&0\\ 0&1&0&(1-\delta)^{1/4}\\ 1&0&-1/(1-\delta)^{1/4}&0\\ 0&1&0&-(1-\delta)^{1/4}\end{array}\right),$ (B31)
@@
Appendix C Cosmic-ray collisions
To evaluate shielding performance against cosmic-ray-induced test mass excitations, we performe Monte Carlo simulations using Geant4 [30]. The test mass is modeled as a cylindrical Pt-Au alloy (27% Pt and 73% Au by mass), located in the center of a spherical multilayer shield. We consider a mass value of 18 mg for the test mass. The test mass is modeled with radius $R=L\sqrt{2/\pi}$ and height $h=L$, which yields $\Lambda=2$ for $m/L^{3}=\rho\Lambda$.
@@ Appendix C Cosmic-ray collisions
We record the momentum kicks delivered to the test mass by cosmic-ray collisions. To realize GIE, the number of phonons generated by collisions must remain below at most one, throughout the GIE generation time $2\pi/(\delta\Omega)$. In practice, if a phonon excitation does occur, measurements are paused until the test mass has returned to its initial state. By increasing the feedback cooling gain $g_{0}$ to unity after each excitation, the effective quality factor $Q$ is reduced to 1. Consequently, the relaxation time scale to the initial state is approximately $2\pi/\Omega$, and the initialization time is weighted according to the noise magnitude. By performing the simulation described above, we find that the duty cycle, the fraction of time the measurement process is active within a given period, reaches 56% with 1.26 m of iron shielding. The average total number of large events during a measurement period of $2\pi/(\delta\Omega)$ is 0.895. The top panel of Fig. 5 shows the simulation result.
Figure 5: Simulation of test mass vibrations caused by cosmic-ray collisions, for a mass of 18 mg and $\delta=0.48$. Top panel: The time series data for the first $10^{5}$ seconds. The red segments (on a gray background) indicate regions excluded by data processing. The black dots indicate collision events with cosmic rays. Bottom panel: The averaged power spectral density obtained exclusively from the blue segments that were not excluded during data processing and whose durations exceed $2\times 2\pi/(\delta\Omega)$. It is comparable to the thermal noise level (dashed red).
In terms of the acceleration noise level due to cosmic-ray collisions, the required value for the amplitude spectral density (ASD) is calculated by the fluctuation-dissipation theorem [48]:
$\sqrt{S_{a}}$ $\simeq\frac{3.1\times 10^{-18}\ {\rm m}}{\rm{s}^{2}\sqrt{\text{Hz}}}\sqrt{\left(\frac{18~{}\text{mg}}{\rm{m}}\right)\left(\frac{\Gamma T/2\pi}{10^{-18}~{}\text{Hz}\cdot\text{K}}\right)}.$ (C32)
The requirement is met with 1.26 meters of iron shielding, as shown in the bottom panel of Fig. 5, where the acceleration noise is normalized by the zero-point amplitude. Although a total thickness of 1.5 m may not be impossible, it should be made thinner. It can be achieved through optimization of the shielding materials, implementation of active shielding with superconducting coils, and reduction of the test mass size.
Finally, in terms of electron charging, the rate due to secondary electrons generated by primary cosmic-ray protons is $7\times 10^{-5}\,\mathrm{e/s}$ (2200 electrons per year), which is approximately six orders of magnitude lower than the value measured by LPF [45]. The ratio of the gravitational potential energy $|\phi_{\rm Grav.}|=Gm^{2}/L$ and the Coulomb potential energy $|\phi_{\rm Coul.}|=N^{2}e^{2}/(4\pi\varepsilon_{0}L)$ is
$\biggl{|}{\phi_{\rm Grav.}\over\phi_{\rm Coul.}}\biggr{|}=3\times 10^{9}\left({m\over 100\ {\rm mg}}\right)^{2}{1\over N^{2}},$ (C33)
where $N$ is the number of electrons. This means that the effect of the Coulomb potential is negligible as long as $N<5\times 10^{4}$ for $m=100$ milligram objects.
Appendix D Probability density functions
We compute the probability density functions of $\Sigma-\det(\bm{\mathcal{V}})$, $\bm{V}_{+}$, and $\bm{V}_{-}$, with four integration (measurement) times: $10^{5}$ s, $10^{6}$ s, $5\times 10^{6}$ s , and $10^{7}$ s. In the top and middle panels of Fig. 6, we show the probability density functions of $\Sigma-\det(\bm{\mathcal{V}})$ and of $\bm{V}_{+}$ and $\bm{V}_{-}$, respectively, for three integration times: $10^{6}$ s (blue), $5\times 10^{6}$ s (green), and $10^{7}$ s (red). The asterisks indicate the respective theoretical values. Compared to the differential mode, the conditional variance of the common mode exhibits a slight deviation from the theoretical value; this asymmetry arises because the laser power is equal in both modes, and the higher resonance frequency of the common mode makes it more susceptible to the influence of shot noise. In the lower panel of Fig. 6, circle and star symbols represent $P_{\mathrm{obs}}^{\mathrm{GIE}}$ for $\delta=0.28$ and $\delta=0.48$, respectively. The blue and red lines represent the cases where the cumulative probability increases in proportion to the square root of the integration time.

### v2

@@ Appendix A State space model
$\hat{p}(t)$ $=\omega_{\mathrm{fb}}\int^{t}\!dt^{\prime}\,e^{-\omega_{\mathrm{fb}}(t-t^{\prime})}\frac{\dot{Y}(t^{\prime})}{-\alpha\Omega}$ (A11)
$=-\,\frac{\omega_{\mathrm{fb}}}{\alpha\Omega}\,Y(t)+\frac{\omega_{\mathrm{fb}}^{2}}{\alpha\Omega}\int^{t}\!dt^{\prime}\,e^{-\omega_{\mathrm{fb}}(t-t^{\prime})}Y(t^{\prime}),$
@@
$r(t)=\frac{\omega_{\mathrm{fb}}}{\alpha}\int^{t}\!dt^{\prime}\,e^{-\omega_{\mathrm{fb}}(t-t^{\prime})}\,Y(t^{\prime}),$ (A12)
@@
Thus, $\gamma_{\pm}$ represent the full width at half-maximum (FWHM). To demonstrate the GIE, the condition $\gamma_{+}>2\Gamma n_{+}$ represents an approximate threshold, as depicted by the red curve in Fig. 2 of the main text.
Appendix B Form factor
For two point masses $m$ separated by a distance $L$, the interaction Hamiltonian expanded to the second order in the differential displacement $q$ is:
$H_{\text{int}}$ $=\frac{m\Omega^{2}}{2}q^{2}-\frac{Gm^{2}}{L+q}$
$\approx\frac{m\Omega^{2}}{2}q^{2}-\left(\frac{Gm^{2}}{L}-\frac{Gm^{2}}{L^{2}}q+\frac{Gm^{2}}{L^{3}}q^{2}\right)$ (B19)
The term proportional to $q^{2}$ modifies the effective spring constant of the system. We can express this modification using the dimensionless gravitational coupling parameter $\delta$:
$H_{\text{int}}\approx\frac{m\Omega^{2}(1-\delta)}{2}q^{2},\quad\text{where}\quad\delta=\frac{4Gm}{\Omega^{2}L^{3}}$ (B20)
for the point-mass case.
However, our experimental setup uses mirrors of finite cylindrical size. To account for the deviation from the point-mass approximation, we introduce a form factor $\Lambda$. This factor modifies the gravitational potential and, consequently, the coupling parameter $\delta$:
$H_{\text{int, finite}}=\frac{m\Omega^{2}}{2}q^{2}-\Lambda\frac{Gm^{2}}{L+q}$ (B21)
This leads to a corrected coupling parameter:
$\delta=\Lambda\frac{4Gm}{\Omega^{2}L^{3}}.$ (B22)
The form factor $\Lambda$ represents the ratio of the gravitational force gradient for finite-sized cylinders to that for point masses. More precisely, it is defined as the ratio of the second derivatives of the gravitational potential energy with respect to the center-of-mass separation $L$:
$\Lambda=\frac{\frac{\partial^{2}U_{\text{cylinder}}}{\partial L^{2}}}{\frac{\partial^{2}U_{\text{point-mass}}}{\partial L^{2}}}=\frac{\frac{\partial^{2}U_{\text{cylinder}}}{\partial L^{2}}}{\frac{2Gm^{2}}{L^{3}}}$ (B23)
where $U_{\text{cylinder}}$ is the gravitational potential energy between two identical, coaxial cylinders, each with mass $m$, radius $R$, and height $h$. Their centers of mass are separated by a distance $L$, resulting in a face-to-face separation of d=Lhd=L-h.
The calculation of $U_{\text{cylinder}}$ involves integrating the Newtonian potential over the volumes of both cylinders. While the full derivation is extensive, the resulting expression for $\Lambda$ can be formulated as a function of two dimensionless geometric ratios: the radius-to-height ratio ($R/h$) and the center-of-mass–to-height ratio ($L/h$).
Based on established methods for calculating gravitational interactions between cylindrical bodies, the form factor $\Lambda$ can be expressed by the following integral:
$\Lambda=8\frac{(d+h)^{3}}{R^{2}h}\int_{0}^{\infty}\left[\frac{J_{1}\left(\frac{R}{h}s\right)}{s}\right]^{2}e^{-\frac{d}{h}s}\sinh^{2}\left(\frac{s}{2}\right)ds$ (B24)
Here, $J_{1}(x)$ is the Bessel function of the first kind of order one.
Due to the complexity of the integrand in Eq. (B24), we calculated the form factor $\Lambda$ numerically. The results of this numerical calculation are shown in Figure 5. The plot illustrates how $\Lambda$ varies as a function of the cylinder geometry. As expected, $\Lambda$ approaches unity for large separations ($L/h\gg 1$), where the finite-size effects become negligible and the cylinders behave like point masses. For closer separations and larger radii, $\Lambda$ deviates significantly from 1, highlighting the necessity of this correction for an accurate analysis of the system.
Figure 5: The calculated form factor $\Lambda$ as a function of the dimensionless geometric parameters $L/h$ and $R/h$. The value of $\Lambda$ converges to 1 for large separations, corresponding to the point-mass approximation. The red cross mark corresponds to the parameters selected in this study.
Appendix C Covariance matrix
@@
$\bm{\tilde{V}}_{\pm}=\left(\begin{array}[]{cc}V^{\pm}_{qq}&V^{\pm}_{qp}\\ V^{\pm}_{qp}&V^{\pm}_{pp}\end{array}\right).$ (C27)
@@
$\bm{\mathcal{V}}\equiv\Bigl(\begin{array}[]{cc}\bm{\mathcal{V}}_{1}&\bm{\mathcal{V}}_{12}\\ \bm{\mathcal{V}}_{12}&\bm{\mathcal{V}}_{2}\end{array}\Bigr)=\bm{S}\Bigl(\begin{array}[]{cc}\bm{\tilde{V}}_{+}&0\\ 0&\bm{\tilde{V}}_{-}\end{array}\Bigr)\bm{S}^{\text{T}},$~~ (C32)
$\bm{S}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cccc}1&0&1/(1-\delta)^{1/4}&0\\ 0&1&0&(1-\delta)^{1/4}\\ 1&0&-1/(1-\delta)^{1/4}&0\\ 0&1&0&-(1-\delta)^{1/4}\end{array}\right),$ (C37)
@@
Appendix D Cosmic-ray collisions
To evaluate cosmic-ray–induced excitations of the test mass, we perform Monte Carlo simulations using Geant4-11.3.1 [30] and Root 6.34.08 [30]. The test mass is modeled as a cylindrical Pt–Au alloy (27% Pt and 73% Au by mass) with a total mass of 1 kg. The required amplitude spectral density (ASD) of the acceleration noise is calculated using the fluctuation–dissipation theorem. [43]:
$\sqrt{S_{a}^{(\rm req)}}$ $\simeq\frac{5.9\times 10^{-21}\ {\rm m}}{\rm{s}^{2}\sqrt{\text{Hz}}}\sqrt{\left(\frac{1~\text{kg}}{\rm{m}}\right)\left(\frac{\Gamma T/2\pi}{2\times 10^{-19}~\text{Hz}\cdot\text{K}}\right)}.$ (D38)
In the simulation, a test mass is placed at the center of the sphere. Protons are injected from the spherical surface with directions cosine-biased relative to the local inward normal. Their energies follow the known galactic cosmic-ray flux distribution [31] [32], with energies ranging from $10^{2}$ to $10^{5}$ MeV.
We record the momentum kicks delivered to the test mass by cosmic‐ray collisions. To realize GIE, the number of phonons generated by collisions must remain below at most one, throughout the GIE generation time $2\pi/(\delta\Omega)=4400$ s for $\Omega/(2\pi)=0.25$ mHz and $\delta=0.91$. In practice, if a phonon excitation does occur, measurements are paused until the test mass has returned to its initial state. By increasing the feedback cooling gain $g_{0}$ to unity after each excitation, the effective quality factor $Q$ is reduced to 1. Consequently, the relaxation time scale to the initial state is approximately $2\pi/\Omega$, and the initialization time is weighted according to the noise magnitude. By performing the simulation described above (see the top panel of Fig. 6), we find that the requirement is satisfied with a duty cycle of 12%, defined as the fraction of time the measurement process is active within a given period (see the bottom panel of Fig. 6). The average total number of large events during a measurement period of $2\pi/(\delta\Omega)$ is 1.9.
Figure 6: Simulation of test mass vibrations caused by cosmic-ray collisions, for a mass of 1 kg and $\delta=0.91$. Top panel: The time series data for the first 100,000 seconds of the 1.1-million-second simulation. The red segments (on a gray background) indicate regions excluded by data processing. The black dots indicate collision events with cosmic rays. Bottom panel: The averaged power spectral density obtained exclusively from the blue segments that were not excluded during data processing and whose durations exceed $2\pi/(\delta\Omega)$. It is comparable to the thermal noise level (dashed red).
Appendix E Probability density functions
We compute the probability density functions of $\Sigma-\det(\bm{\mathcal{V}})$, $\bm{V}_{+}$, and $\bm{V}_{-}$ for two distances: $L=1.25h\ (\delta=0.91)$ and $L=4h/3\ (\delta=0.75)$. In the top and middle panels of Fig. 7, we show the probability density functions of $\Sigma-\det(\bm{\mathcal{V}})$ and of $\bm{V}_{+}$ and $\bm{V}_{-}$, respectively, for three integration times: $5\times 10^{5}$ s (blue), $2\times 10^{6}$ s (green), and $3\times 10^{6}$ s (red), fot the $\delta=0.91$ case. The asterisks indicate the respective theoretical values.
Compared to the differential mode, the conditional variance of the common mode exhibits a slight deviation from the theoretical value; this asymmetry arises because the laser power is equal in both modes, and the higher resonance frequency of the common mode makes it more susceptible to the influence of shot noise. In the lower panel of Fig. 7, circle and star symbols represent $P_{\mathrm{obs}}^{\mathrm{GIE}}$ for $\delta=0.75$ and $\delta=0.91$, respectively. The blue and red lines represent the cases where the cumulative probability increases in proportion to the square root of the integration time.
Appendix F Passive Shielding for a Common-Mode Inertial Reference
F-1. The Challenge of Common-Mode Acceleration Noise
As discussed in the main text, our proposed experiment requires simultaneous measurements of both the differential and common-mode motions of two test masses (TMs). While the differential measurement is largely insensitive to the spacecraft’s motion, the common-mode measurement is fundamentally limited by the inertial stability of the spacecraft itself. Consequently, any acceleration of the Science Module (SM) that houses the TMs manifests as common-mode acceleration noise.
@@ F-1. The Challenge of Common-Mode Acceleration Noise
To enable GIE detection, the acceleration-noise amplitude spectral density of the 1-kg test masses must satisfy
$\sqrt{S_{a,\mathrm{TM}}^{(\mathrm{req})}}<5.9\times 10^{-21}~\mathrm{m/s^{2}/\sqrt{Hz}},$ (F39)
evaluated at the resonance $f=0.25~\mathrm{mHz}$. This level is several orders of magnitude below what conventional active drag-free control can achieve, motivating a passively shielded configuration.
F-2. Shielded Formation Concept
Two modules fly in loose formation:
  1. 1.
    Science Module (SM): The 700-kg spacecraft contains the test masses (TMs), auxiliary masses, interferometer, vacuum systems, and refrigeration systems. Reference cross-sectional area $A_{\mathrm{SM}}=3.5~\mathrm{m^{2}}$, giving an equivalent diameter $D_{\mathrm{SM}}=2.11~\mathrm{m}$.
  2. 2.
    Shield Module (ShM): A tungsten-alloy disk positioned between the Sun and the SM at separation $L=275~\mathrm{m}$. To maintain a full umbra over the SM,
    $d_{\mathrm{ShM}}\geq D_{\mathrm{SM}}+L\,\theta_{\odot},\ \theta_{\odot}=9.3\times 10^{-3}~\mathrm{rad}.$ (F40)
    Hence $d_{\mathrm{ShM}}=4.67~\mathrm{m}$ ensures complete shadowing. For thickness $t_{\mathrm{ShM}}=0.02~\mathrm{m}$ and density $\rho=19{,}300~\mathrm{kg/m^{3}}$,
    $m_{\mathrm{ShM}}=\pi(d/2)^{2}t\rho\approx 6.6\times 10^{3}~\mathrm{kg}.$ (F41)
    The Sun-facing surface, with a solar absorptance of $\simeq 0.1$ and an emissivity of $\simeq 0.9$, is passively maintained at approximately $T_{\mathrm{ShM}}\approx 227~\mathrm{K}$.
F-2. Transfer Function Between SM and TM
The apparent common-mode acceleration of each TM, referred to the optical bench on the SM, follows the SM acceleration through the TM–SM coupling. Modeling this coupling as a lightly damped harmonic link with natural frequency $\Omega$ (set by auxiliary masses) and damping ratio $\zeta=1/(2Q)$ gives
$H(f)=\frac{(2\pi f)^{2}}{(2\pi f)^{2}-\Omega^{2}+i\,2\zeta\,\Omega(2\pi f)}.$ (F42)
At the resonance frequency $f_{0}=\Omega/2\pi$, the magnitude increases to $|H(f_{0})|=1/(2\zeta)=Q$. This transfer function multiplies SM-originated acceleration noises when referred to the TM readout. The quality factor $Q$ thus determines how strongly SM motion is transmitted to the apparent TM acceleration.
F-3. Residual Acceleration Noise (SM Level)
The residual acceleration noise acting on the SM arises from two mechanisms: (i) gravitational coupling to ShM motion driven by solar-radiation-pressure (SRP) fluctuations, and (ii) fluctuating thermal radiation pressure. We adopt $\sqrt{S_{T}}=0.1~\mathrm{K/\sqrt{Hz}}$ and $\mathcal{G}_{\mathrm{sup}}=10^{-6}$ for geometric suppression.
1. Gravitational Coupling from SRP Fluctuations
The time-varying Newtonian field produced by fluctuations of the Shield Module (ShM) acts on both the Science Module (SM) and the TMs. Because the ShM–SM separation ($L$) is hundreds of meters whereas the internal TM–SM spacing ($s$) is only on the order of tens of centimeters, the gravitational acceleration from the ShM is almost identical at the two locations. As a result, the SM and TMs experience nearly the same instantaneous acceleration, so that in the relative (TM–SM) coordinate this common term cancels to first order. The residual signal arises only from the gradient of the ShM’s field, that is, from the small difference in gravitational pull across the TM–SM baseline.
Expanding $g(r)=Gm_{\rm Sh}/r^{2}$ about $r=L$ and retaining the leading gradient term $({\rm d}g/{\rm d}r)_{L}\,s=-2Gm_{\rm Sh}s/L^{3}$, a small displacement $x_{\rm Sh}$ of the ShM produces a differential acceleration between the TM and SM of amplitude $\delta a_{\rm rel}\simeq 6Gm_{\rm Sh}s\,x_{\rm Sh}/L^{4}$. Relating the ShM displacement noise to its force noise through $x_{\rm Sh}=\sqrt{S_{F,{\rm ShM}}}/(m_{\rm Sh}\omega^{2})$ then gives
$\sqrt{S_{a,{\rm TM}\!-\!{\rm SM}}^{\rm(ShM\ grav.)}}\simeq\frac{6\,G\,s}{L^{4}}\,\frac{\sqrt{S_{F,{\rm ShM}}}}{\omega^{2}},\qquad\omega=2\pi f.$ (F43)
Numerically this yields $\sim 2\times 10^{-23}~\mathrm{m/s^{2}/\sqrt{Hz}}$ for $f=0.25~\mathrm{mHz}$, $L=275~\mathrm{m}$, $s\simeq 0.5~\mathrm{m}$, and $\sqrt{S_{F,{\rm ShM}}}=1.7~\mathrm{nN/\sqrt{Hz}}$. This coupling path does not include the $H(f)$ gain factor, because the first-order common acceleration of the SM and TM is suppressed by common-mode cancellation, leaving only the much smaller tidal (gravity-gradient) term in the relative readout.
2. Thermal Radiation Pressure
Temperature fluctuations on the Sun-facing surface of the Shield Module lead to variations in its thermal emission, producing a fluctuating radiation pressure on the Science Module. This effect is strongly suppressed by geometry and baffling, but its residual contribution can be estimated by linearizing the Stefan–Boltzmann law with respect to temperature fluctuations $\delta T$. The corresponding acceleration noise on the SM is given by
$\sqrt{S_{a,\mathrm{SM}}^{\mathrm{th}}}$ $=\frac{1}{m_{\mathrm{SM}}}\left(\frac{4\,\varepsilon\,\sigma\,T_{\mathrm{ShM}}^{3}}{c}\right)\left(\frac{A_{\mathrm{ShM}}}{4\pi L^{2}}\right)\sqrt{S_{T}}\,G_{\mathrm{sup}}$
$\approx 7\times 10^{-23}~\mathrm{m/s^{2}/\sqrt{Hz}}.$ (F44)
This SM-only path does map to the TM readout with the transfer gain $|H|$.
This expression assumes isotropic thermal emission from the shield surface, with $\mathcal{G}_{\mathrm{sup}}$ representing the net geometric suppression factor due to multi-stage baffling and low-emissivity internal coatings.
F-4. Summary and Optimum $Q$
The TM-side total at $f=0.25~\mathrm{mHz}$ combines (i) the gravity-gradient residue from the ShM (no $H$ factor) and (ii) the SM-only thermal path scaled by $|H|=Q$:
$\sqrt{S_{a,\mathrm{TM}}^{\mathrm{total}}}=\sqrt{\left(\sqrt{S_{a,{\rm TM}\!-\!{\rm SM}}^{\rm(ShM\ grav.)}}\right)^{2}+\left(Q\,\sqrt{S_{a,\mathrm{SM}}^{\mathrm{th}}}\right)^{2}}\,.$ (F45)
With the present numbers $\sqrt{S_{a,{\rm TM}\!-\!{\rm SM}}}^{\rm(ShM\ grav.)}\!\approx 2\times 10^{-23}$ and $\sqrt{S_{a,\mathrm{SM}}^{\mathrm{th}}}\!\approx 7\times 10^{-23}$, the requirement $\sqrt{S_{a,\mathrm{TM}}^{(\mathrm{req})}}=5.9\times 10^{-21}$ is satisfied provided
[…]
Thus, the design easily meets the requirement for any practical $Q$ up to $\mathcal{O}(10^{2})$ under the present thermal assumptions. If, for conservatism, one ignores common-mode cancellation and upper-bounds the SM-only path by the absolute SM acceleration level $\sim 4.4\times 10^{-21}~\mathrm{m/s^{2}/\sqrt{Hz}}$, the resulting limit is $Q\lesssim 1.3$, which remains achievable.
The present configuration — a 4.7 m-diameter, 6.6 t tungsten shield located at L=275 m — therefore satisfies the common-mode inertial-stability requirement. Future work will further refine the design by (i) quantifying the geometric and control tolerances needed to maintain full umbra and formation stability, (ii) developing a detailed force-noise and thermal-fluctuation budget for both the shield and the spacecraft, and (iii) assessing the feasibility of power transfer and attitude control under realistic operational conditions.